OpenStudy (anonymous):

How do u make a perpendicular through y=4x+2 and goes through (8,-2)

OpenStudy (anonymous):

RIGHT NOW in you math problem it only deals with SLOPE so... 

OpenStudy (anonymous):

that is an example

OpenStudy (ranga):

slopes of perpendicular lines are negative reciprocal of each other.

OpenStudy (anonymous):

Thanks

OpenStudy (ranga):

y = 4x + 2 The slope of this line is 4 The slope of the perpendicular line is -1/4 y = mx + b y = -1/4x + b It passes through (8, -2) -2 = -1/4*8 + b -2 = -2 + b b = 0 So the perpendicular line passing through (8,-2) is: y = -1/4x